# Accelerated Gradient Methods with Biased Gradient Estimates: Risk Sensitivity, High-Probability Guarantees, and Large Deviation Bounds

### v1

@@ 3. Preliminaries
$y_{k}$ $=$ $x_{k}+\nu(x_{k}-x_{k-1})$ (3.1b)
starting from an arbitrary initial point $x_{0}\in\mathbb{R}^{d}$ with the convention that $x_{-1}=x_{0}$. GMM admits three non-negative parameters $\alpha,\beta$ and $\nu$, where $\alpha>0$ is the step size, and $\beta$ and $\nu$ are called “momentum parameters". This method generalizes a number of momentum-averaging-based first-order algorithms. If $\nu=0$, GMM is equivalent to the Heavy-Ball method (HB) of Polyak [8]. If we choose $\beta=\nu$, GMM recovers Nesterov’s accelerated gradient (NAG) method [9]. On the other hand, when $\beta=\nu=0$, GMM reduces to the gradient descent (GD) method. Another method, the triple momentum method (TMM) corresponds to the parameter choice $\alpha=\tfrac{1+\bar{\rho}}{L}$, $\beta=\tfrac{\bar{\rho}^{2}}{2-\bar{\rho}}$, $\nu=\tfrac{\bar{\rho}^{2}}{(1+\bar{\rho})(2-\bar{\rho})}$, where $\bar{\rho}=1-1/\sqrt{\kappa}$ [10].
A growing body of work has employed control-theoretic tools to analyze the robustness of momentum-based optimization algorithms to noise [53, 57, 12, 94, 95, 55]. To the best of our knowledge, our work is the first to investigate the risk sensitivity of the dynamical systems associated with GMMs, with the aim of characterizing the trade-offs between their sensitivity to gradient noise and their convergence rates. We study the performance of GMM algorithms subject to additive gradient noise, i.e., at each iteration $k\geq 0$, instead of employing the actual gradient $\nabla f(y_{k})$, suppose we only have access to its noisy estimate:
@@
where $w_{k+1}\in\mathbb{R}^{d}$ is the additive gradient noise at $y_{k}$. Assumptions on the exact nature of this noise will be provided later. Here, we present the dynamical system reformulation of (3.1) as a discrete-time dynamical system:
@@
$y_{k}=C\xi_{k},\quad u_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k+1}}=\tilde{\nabla}f(y_{k},w_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k+1}})=\nabla f(y_{k})+w_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k+1}},$ (3.3b)
@@
$A=\tilde{A}\otimes\mathsf{I}_{d},\quad B=\tilde{B}\otimes\mathsf{I}_{d},\quad C=\tilde{C}\otimes\mathsf{I}_{d},$ (3.4)
@@
Assumption 3.1 encompasses a variety of noise models relevant to practical scenarios. The centered (unbiased) stochastic noise is commonly encountered in differential privacy, where i.i.d. centered noise is added to gradients to protect user data [18]. A similar situation arises in stochastic optimization when the objective takes the form $f(x)\coloneqq\mathbb{E}_{\xi}[g(x,\xi)]$, and the expectation is approximated via empirical averaging over mini-batches. Under moderate batch sizes, the Central Limit Theorem justifies modeling the resulting gradient noise as Gaussian, a behavior also observed empirically [22, 23]. Consequently, Assumption 3.1 is often satisfied in practice. Light-tail conditions such as this are standard in the analysis of stochastic optimization algorithms, particularly for deriving tail bounds and high-probability guarantees. On the other hand, it is possible that the gradient noise is deterministic, i.e., there is no stochastic element the noise. This situation is typical in incremental gradient methods or in settings where gradients are approximated by solving subproblems inexactly [13, 99, 98]. Our noise model also captures deterministic corruption of gradients at each iteration, which may arise due to low-precision floating-point arithmetic [100].
@@ 3.1. Risk-sensitive cost and the risk-sensitive index
provided that there exists some $\bar{\theta}>0$ such that $R_{K}(\theta)$ is finite for $\theta\in(0,\bar{\theta})$. The parameter $\theta$ is called the risk-sensitivity parameter [91, 62]. The term $S_{K}$ is a running sum of suboptimality over $K$ iterations; therefore, we see from (3.11) that when $\theta=0$, the risk-sensitive cost is a measure of average performance $S_{K}/(K+1)$ in expectation. For $\theta>0$, as $\theta$ gets larger, the risk-sensitive cost $R_{K}(\theta)$ penalizes large values of $S_{K}$ more and more due to the exponential term in (3.10) and $R_{K}$ will be more sensitive to large values of suboptimality. Therefore, in analogy with a risk-averse financial investor that tries to avoid large losses, the case $\theta>0$ is called the “risk-averse case" (see e.g. [61, 91]). This can also be readily seen from the Taylor expansion of (3.10) around $\theta=0$, i.e.,
@@
provided that the limit exists. Risk-sensitive cost is well-studied in the robust control literature and is used to quantify the robustness of dynamical systems to stochastic noise [62, 102, 103, 104]. In some of our results, we will also consider the $\theta<0$ case, i.e., risk-seeking scenario.
3.2. $H_{\infty}$ norm.
@@
$\|z_{k}\|^{2}=\|F(\xi_{k})\|^{2}=f(x_{k})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}},$ (3.13)
for all $k\geq 0$, then the cumulative cost $S_{K}=\sum_{k=0}^{K}\big{(}f(x_{k})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}\big{)},$ can be expressed in terms of the $\ell_{2}$ norm of the sequence $\{z_{k}\}$, namely, $S_{K}=\sum_{k=0}^{K}\|z_{k}\|^{2}$ –for instance, the condition (3.13) is satisfied when the weighting function in (3.9) is chosen as $\overline{w}(z)=\sqrt{z}$. In this case, when the noise $w_{k+1}$ is deterministic for every $k\geq 0$ and is square-summable; from the limit (3.11), we see that $R_{K}(0)=\frac{S_{K}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K+1}}$ is also deterministic and corresponds to the average suboptimality. Given the decaying nature of square-summable noise, a natural relevant question would be whether the deterministic limit $S_{\infty}:=\lim_{K\to\infty}S_{K}=\sum_{k=0}^{\infty}\|z_{k}\|^{2}$, which captures the cumulative suboptimality over the iterations, exists –with the convention that the limit is $\infty$ if $S_{K}$ is not bounded.
@@
The $H_{\infty}$-norm quantifies the worst-case deterministic noise amplification of the system in (3.8), i.e., it quantifies how the output behaves in response to the noise input $\{w_{k+1}\}_{k\geq 0}$. A smaller $H_{\infty}$ value indicates a more robust system, as the output is less sensitive to worst-case square-summable input noise; consequently, the cumulative suboptimality over the iterations, i.e., $\sum_{k\geq 0}\big{(}f(x_{k})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}\big{)}$, will also be smaller for the same noise level. For GMM, $H_{\infty}$-norm depends on the algorithm parameters $(\alpha,\beta,\nu)$ and the problem parameters $L$ and $\mu$, i.e., smoothness and convexity constants, respectively. For strongly convex quadratics, it is known that $H_{\infty}$-norm can be computed exactly. Indeed, the following result is a direct consequence of [57, Theorem 4.1] and its proof is based on [57, eqn. (6), Lemma A.1].
@@
Finally, it always hold that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}H_{\infty}}\geq\frac{1}{\sqrt{2\mu}}$.
For general strongly convex functions that are not quadratic, exact analytical expressions are typically unavailable. Nevertheless, tight upper bounds have been derived in [57] for GD, NAG, and more generally for GMM, by verifying the feasibility of certain $4\times 4$ matrix inequalities. For completeness, we present these bounds in theorem˜A.1 in the appendix. In the context of stochastic dynamical systems driven by Gaussian noise, it is known that risk sensitivity is closely related to the $H_{\infty}$ norm: specifically, the largest value of $\theta>0$ for which the risk-sensitive index $R(\theta)$ remains finite depends on the system’s $H_{\infty}$ norm [103, 61]. To the best of our knowledge, however, such connections have not been explored in the setting of optimization algorithms driven by noise models such as those described in ˜3.1. To investigate this connection further and gain insight via precise analytical formulas, the next section focuses on the special case of strongly convex quadratic objectives subject to unbiased (centered) i.i.d. Gaussian noise. This serves as a foundation before extending the analysis to general smooth strongly convex functions under the broader noise model of ˜3.1.
@@ 4. Exact analysis of $R(\theta)$ for strongly convex quadratics
and, from (4.3), it can be seen that $z_{k}$ satisfies the identity (3.13). Therefore, based on eq.˜3.14, we can talk about the $H_{\infty}$-norm of the GMM system and the $H_{\infty}$ formula from theorem˜3.4 is applicable. The $H_{\infty}$ norm characterizes robustness in the hypothetical setting where the noise is fully deterministic. In ˜3.1, we consider more general noise models; however, if the noise specified therein were further restricted to be deterministic and square-summable, the $H_{\infty}$ norm would quantify robustness with respect to the worst-case impact of such noise.
The GMM dynamics, as defined by (4.5)–(4.7), constitute a linear dynamical system identified by the system matrices $(A_{Q},B,T)$. In this context, in the special case when the noise sequence $\{w_{k+1}\}_{k\geq 0}$ consists of i.i.d. multivariate Gaussian vectors such that $w_{k+1}\sim\mathcal{N}(0,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\sigma^{2}}{d}{\mathsf{I}_{d}}}})$ for some $\sigma>0$,it is known that the risk-sensitive index $R(\theta)$ is finite if and only if $\sqrt{\theta}H_{\infty}<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{d}}$ [61]; indeed, a particular form of a discrete-time algebraic Riccati equation (DARE), i.e.,
@@
admits a positive semi-definite solution $X\succeq 0$ if and only if $\gamma>H_{\infty}$ [109, Corollary 2.1, Prop. 3.7] –thus, $\sqrt{\theta}H_{\infty}<\sqrt{d}$ holds if and only if (4.8) has a positive semi-definite solution for the parameter choice $\gamma=\frac{1}{\sqrt{\theta/d}}$ (see [93, Lemma 3.4] and [61]). To compute a solution $X\in\mathbb{R}^{2d\times 2d}$ for the DARE above, a standard method involves tackling a $4d\times 4d$ generalized eigenvalue problem. More specifically, it necessitates the calculation of the generalized eigenvalues and eigenvectors of the following $4d\times 4d$ matrix pencil:
@@
[…]
@@
In the setting of theorem˜4.1, given step size $\alpha\in(0,\frac{2}{L})$, consider the gradient descent method, i.e., $\beta=\nu=0$, the risk-sensitive index for the risk parameter $\theta>0$ is given by
$R(\theta)=\begin{cases}-\frac{\sigma^{2}}{\theta}\sum_{i=1}^{d}\log\left(1-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\theta}{d}}\alpha^{2}\tilde{a}_{i}\right)&\mbox{if}\quad\sqrt{\theta}H_{\infty}<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{d}},\\ \infty&\mbox{if}\quad\sqrt{\theta}H_{\infty}\geq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{d}},\end{cases}$ (4.12)
@@
$H_{\infty}=\begin{cases}\frac{1}{\sqrt{2\mu}}&\text{if}\quad 0<\alpha\leq\frac{2}{L+\sqrt{L\mu}},\\ \frac{\alpha\sqrt{L}}{\sqrt{2(2-\alpha L)}}&\text{if}\quad\frac{2}{L+\sqrt{L\mu}}<\alpha<\frac{2}{L},\end{cases}$ (4.13)

### v2

@@ 3. Preliminaries
$y_{k}$ $=$ $x_{k}+\nu(x_{k}-x_{k-1}),$ (3.1b)
starting from an arbitrary initial point $x_{0}\in\mathbb{R}^{d}$ with the convention that $x_{-1}=x_{0}$. GMM admits three non-negative parameters $\alpha,\beta$ and $\nu$, where $\alpha>0$ is the step size, and $\beta$ and $\nu$ are called “momentum parameters". This method generalizes a number of momentum-averaging-based first-order algorithms. If $\nu=0$, GMM is equivalent to the heavy-ball method (HB) of Polyak [8]. If we choose $\beta=\nu$, GMM recovers Nesterov’s accelerated gradient (NAG) method [9]. On the other hand, when $\beta=\nu=0$, GMM reduces to the gradient descent (GD) method. Finally, the triple momentum method (TMM) corresponds to the parameter choice $\alpha=\tfrac{1+\bar{\rho}}{L}$, $\beta=\tfrac{\bar{\rho}^{2}}{2-\bar{\rho}}$, $\nu=\tfrac{\bar{\rho}^{2}}{(1+\bar{\rho})(2-\bar{\rho})}$, where $\bar{\rho}=1-1/\sqrt{\kappa}$ [10].
A growing body of work has employed control-theoretic tools to analyze the robustness of momentum-based optimization algorithms to noise [53, 57, 12, 95, 96, 55]. To the best of our knowledge, our work is the first to investigate the risk sensitivity of the dynamical systems associated with GMMs, with the aim of characterizing the trade-offs between their sensitivity to gradient noise and their convergence rates. [^1] Throughout this paper, we use the terms gradient noise and gradient errors interchangeably. We study the performance of GMM algorithms subject to additive gradient noise, i.e., at each iteration $k\geq 0$, instead of employing the actual gradient $\nabla f(y_{k})$, suppose we only have access to its noisy estimate:
@@
where $w_{k+1}\in\mathbb{R}^{d}$ is the additive gradient noise at $y_{k}$. Assumptions on the exact nature of this noise will be provided later. We start with presenting the dynamical system reformulation of (3.1) as a discrete-time dynamical system:
@@
$y_{k}=C\xi_{k},\qquad u_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k+1}}=\tilde{\nabla}f(y_{k},w_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k+1}})=\nabla f(y_{k})+w_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k+1}},$ (3.3b)
@@
$A=\tilde{A}\otimes\mathsf{I}_{d},\qquad B=\tilde{B}\otimes\mathsf{I}_{d},\qquad C=\tilde{C}\otimes\mathsf{I}_{d},$ (3.4)
@@
Assumption 3.1 encompasses a variety of noise models relevant to practical scenarios. The centered (unbiased) stochastic noise is commonly encountered in differential privacy, where i.i.d. centered noise is added to gradients to protect user data [18]. A similar situation arises in stochastic optimization and statistical learning contexts when the objective takes the form $f(x)\coloneqq\mathbb{E}_{\xi}[\mathcal{L}(x,\xi)]$ for a loss function $\mathcal{L}$, and the expectation is approximated via empirical averaging over mini-batches [2, 1]. Under moderate batch sizes, the Central Limit Theorem justifies modeling the resulting gradient noise as Gaussian, a behavior also observed empirically [22, 23]. Consequently, Assumption 3.1 is often satisfied in practice. Light-tail conditions such as this are standard in the analysis of stochastic optimization algorithms, particularly for deriving tail bounds and high-probability guarantees. On the other hand, it is possible that the gradient noise is deterministic, i.e., it contains no stochastic element. This situation is typical in incremental gradient methods or in settings where gradients are approximated by solving subproblems inexactly [13, 100, 99, 101, 102, 6]. Our noise model also captures deterministic corruption of gradients at each iteration, which may arise due to low-precision floating-point arithmetic [103].
@@ 3.1. Risk-sensitive cost and the risk-sensitive index
provided that there exists some $\bar{\theta}>0$ such that $R_{K}(\theta)$ is finite for $\theta\in(0,\bar{\theta})$. The parameter $\theta$ is called the risk-sensitivity parameter [92, 62]. The term $S_{K}$ is a running sum of suboptimality over $K$ iterations; therefore, we see from (3.11) that when $\theta=0$, the risk-sensitive cost is a measure of average performance $S_{K}/(K+1)$ in expectation. For $\theta>0$, as $\theta$ gets larger, the risk-sensitive cost $R_{K}(\theta)$ penalizes large values of $S_{K}$ more and more due to the exponential term in (3.10) and $R_{K}$ will be more sensitive to large values of suboptimality. Therefore, in analogy with a risk-averse financial investor that tries to avoid large losses, the case $\theta>0$ is called the “risk-averse case" (see e.g., [61, 92]). This can also be readily seen from the Taylor expansion of (3.10) around $\theta=0$, i.e.,
@@
provided that the limit exists. Risk-sensitive cost is well-studied in the robust control literature and is used to quantify the robustness of dynamical systems to stochastic noise [62, 105, 106, 107]. In some of our results, we will also consider the $\theta<0$ case, i.e., the risk-seeking scenario.
3.2. $H_{\infty}$-norm.
@@
$\|F(\xi_{k})\|^{2}=f(x_{k})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}},$ (3.13)
for all $k\geq 0$, then the cumulative cost $S_{K}=\sum_{k=0}^{K}\big(f(x_{k})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}\big),$ can be expressed in terms of the $\ell_{2}$ norm of the sequence $\{z_{k}\}$, i.e., since $z_{k}=F(\xi_{k})$ for $k\geq 0$, we have $S_{K}=\sum_{k=0}^{K}\|z_{k}\|^{2}$ –for instance, the condition (3.13) is satisfied when the weighting function in (3.9) is chosen as $\overline{w}(z)=\sqrt{z}$. In this case, when the noise $w_{k+1}$ is deterministic for every $k\geq 0$ and is square-summable; from the limit (3.11), we see that $R_{K}(0)=\frac{S_{K}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K+1}}$ is also deterministic and corresponds to the average suboptimality. Given the decaying nature of square-summable noise, a natural relevant question would be whether the deterministic limit $S_{\infty}:=\lim_{K\to\infty}S_{K}=\sum_{k=0}^{\infty}\|z_{k}\|^{2}$, which captures the cumulative suboptimality over the iterations, exists –with the convention that the limit is $\infty$ if $S_{K}$ is not bounded.
@@
The $H_{\infty}$-norm quantifies the worst-case deterministic noise amplification of the system in (3.8), i.e., it quantifies how the output behaves in response to the noise input $\{w_{k+1}\}_{k\geq 0}$. A smaller $H_{\infty}$ value indicates a more robust system, as the output is less sensitive to worst-case deterministic square-summable input noise; consequently, the cumulative suboptimality over the iterations, i.e., $\sum_{k\geq 0}\big(f(x_{k})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{*}}\big)$, will also be smaller for the same noise level. For GMM, $H_{\infty}$-norm depends on the algorithm parameters $(\alpha,\beta,\nu)$ and the problem parameters $L$ and $\mu$, i.e., smoothness and convexity constants, respectively. For strongly convex quadratics, it is known that $H_{\infty}$-norm can be computed exactly. The following result is a direct consequence of [57, Thm. 4.1] together with [57, eqns. (60)–(61)].
@@
Finally, it always holds that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}H_{\infty}}\geq\frac{1}{\sqrt{2\mu}}$.
For general strongly convex functions that are not quadratic, exact analytical expressions are typically unavailable. Nevertheless, tight upper bounds have been derived in [57] for GD, NAG, and more generally for GMM, by verifying the feasibility of certain $4\times 4$ matrix inequalities. For completeness, we present these bounds in theorem˜A.1 in the appendix. In the context of stochastic dynamical systems driven by Gaussian noise, it is known that risk sensitivity is closely related to the $H_{\infty}$-norm: specifically, the largest value of $\theta>0$ for which the risk-sensitive index $R(\theta)$ remains finite depends on the system’s $H_{\infty}$-norm [106, 61]. To the best of our knowledge, however, such connections have not been explored in the setting of optimization algorithms driven by noise models such as those described in ˜3.1. To investigate this connection further and gain insight via precise analytical formulas, the next section focuses on the special case of strongly convex quadratic objectives subject to unbiased (centered) i.i.d. Gaussian noise. This serves as a foundation before extending the analysis to general smooth strongly convex functions under the broader noise model of ˜3.1.
@@ 4. Exact analysis of $R(\theta)$ for strongly convex quadratics
and, from (4.3), it can be seen that $z_{k}$ satisfies the identity (3.13). Therefore, based on eq.˜3.14, we can talk about the $H_{\infty}$-norm of the GMM system and the $H_{\infty}$ formula from theorem˜3.4 is applicable. The $H_{\infty}$-norm characterizes robustness in the hypothetical setting where the noise is fully deterministic. In ˜3.1, we consider a more general noise model; however, if the noise specified therein were further restricted to be deterministic and square-summable, the $H_{\infty}$-norm would quantify robustness with respect to the worst-case impact of such noise.
The GMM dynamics, as defined by (4.5)–(4.7), constitute a linear dynamical system identified by the system matrices $(A_{Q},B,T)$. In this context, in the special case when the noise sequence $\{w_{k+1}\}_{k\geq 0}$ consists of i.i.d. multivariate Gaussian vectors such that $w_{k+1}\sim\mathcal{N}(0,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\sigma^{2}}{d}{\mathsf{I}_{d}}}})$ for some $\sigma>0$, then the noise variance is $\sigma^{2}$ and ˜3.1 is satisfied. In this case, it is known that the risk-sensitive index $R(\theta)$ is finite if and only if $\sqrt{\theta}H_{\infty}<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{d}}$ [61]; indeed, a particular form of a discrete-time algebraic Riccati equation (DARE), i.e.,
@@
admits a positive semi-definite solution $X\succeq 0$ if and only if $\gamma>H_{\infty}$ [112, Corollary 2.1, Prop. 3.7] –thus, $\sqrt{\theta}H_{\infty}<\sqrt{d}$ holds if and only if (4.8) has a positive semi-definite solution for the parameter choice $\gamma=\frac{1}{\sqrt{\theta/d}}$ (see [94, Lemma 3.4] and [61]). To compute a solution $X\in\mathbb{R}^{2d\times 2d}$ for the DARE above, a standard method involves tackling a $4d\times 4d$ generalized eigenvalue problem. More specifically, it necessitates the calculation of the generalized eigenvalues and eigenvectors of the following $4d\times 4d$ matrix pencil:
@@
where $\gamma=\frac{1}{\sqrt{\theta/d}}$, $\tilde{B}\in\mathbb{R}^{2\times 1}$ is given in (3.5) and
@@
Remark 4.2.
Theorem 4.1 relies on Riccati equations for providing a formula for the risk-sensitive index and applies only when $\theta>0$. Later on, in proposition˜4.4 and corollary˜4.6, we will provide an alternative approach based on the evaluation of an integral that applies to the $\theta\leq 0$ case as well.
@@
Under the premise of theorem˜4.1, given a step size $\alpha\in(0,\frac{2}{L})$, consider the gradient descent method, i.e., $\beta=\nu=0$, the risk-sensitive index corresponding to the risk parameter $\theta>0$ is given by
$R(\theta)=\begin{cases}-\frac{\sigma^{2}}{\theta}\sum_{i=1}^{d}\log\left(1-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{\theta}{d}}\alpha^{2}\tilde{a}_{i}\right)&\mbox{if}\quad\sqrt{\theta}H_{\infty}<{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{d}};\\ \infty&\mbox{if}\quad\sqrt{\theta}H_{\infty}\geq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{d}},\end{cases}$ (4.12)
@@
$H_{\infty}=\begin{cases}\frac{1}{\sqrt{2\mu}}&\text{if}\quad 0<\alpha\leq\frac{2}{L+\sqrt{L\mu}};\\ \frac{\alpha\sqrt{L}}{\sqrt{2(2-\alpha L)}}&\text{if}\quad\frac{2}{L+\sqrt{L\mu}}<\alpha<\frac{2}{L},\end{cases}$ (4.13)